OpenStudy (anonymous):

Part 1: Use complete sentences to describe a real-world scenario that could be represented by the inequality 4x + 3y 36. Part 2: Choose one ordered pair that is a solution to the given inequality and explain what that ordered pair means in the context of your real-world scenario.

OpenStudy (anonymous):

Hm, well you can first change your equation to y= mx + b, perhaps that will be more helpful.

OpenStudy (anonymous):

well i know part 1 is (4/3)x-12

OpenStudy (anonymous):

i just can't figure a real world scenario or part 2

OpenStudy (anonymous):

4x + 3y = 36 3y = -4x + 36 y = -4x/3 + 12

OpenStudy (anonymous):

The slope can represent a rate of anything, since it's negative, it has to represent something with a negative rate.